# Charting the Landscape of African NLP: Mapping Progress and Shaping the Road Ahead

### v2

@@ 4.4 What resources are available?
Given that African languages are low-resource, efforts over the past five years have focused on developing resources—both for individual languages and at a large-scale multilingual level, particularly in the form of datasets. Our analysis shows that 350 papers (about 48%) involve dataset creation, with 106 papers creating dataset via translation. Many large multilingual datasets—such as NLU’s SIB-200 and Taxi-1500—are based on English translations. Similarly, speech datasets like FLEURS, FLEURS-R, and SpeechTaxi rely on translated data, which often lacks cultural relevance and may exhibit translationese (Koppel and Ordan, 2011; Bizzoni et al., 2020), limiting their real-world applicability. Table 1 lists examples of datasets and models, with a more comprehensive list in Appendix E.
@@ 4.5 What are the efficient or low-resource techniques used?
Given the low-resource status of African languages, researchers often rely on specialized techniques. Figure 7 (Appendix D) summarizes those used in the analyzed papers; we highlight five below.
Transfer Learning Our analysis shows that about 44% of the papers used transfer learning, primarily through word embeddings, or by fine-tuning pretrained models or using them as feature extractors. This approach has been effective across many languages, including low-resource ones, largely due to the availability of multilingually pretrained models that support cross-lingual transfer (Pfeiffer et al., 2021). However, a major challenge arises when the target African language is not well represented in the pretrained model or its tokenizer. Common strategies to address this include adaptive pretraining (Pfeiffer et al., 2021; Alabi et al., 2022; Adelani et al., 2024; Meyer et al., 2024), and embedding initialization (Liu et al., 2024; Quinjica and Adelani, 2024; Dobler and de Melo, 2023).
@@
Data Augmentation Data augmentation is another popular low-resource technique used in the analyzed papers covering 6% and it has been sucessfully used for various tasks such as language modeling (Adelani et al., 2024; Singh et al., 2024; Azime et al., 2024), machine translation via back translation (Reid et al., 2021; Adelani et al., 2021b), and other tasks (Zhang et al., 2024b).
@@
Other methods Other techniques including Annotation Projection (AP), and Multi-task learning (MTL) are described in Appendix D.
@@ 4.6 What are the other themes?
Interpretability and analysis (IA) Methods on interpretability and model analysis have gained increasing attention in NLP in recent years (Mosbach et al., 2024). Although few studies focus exclusively on African languages (Chimoto et al., 2024), there are several multilingual IA works on MT (Ahia et al., 2021; Adebara et al., 2022a; Mohammadshahi et al., 2022b), and LLM (Ahia et al., 2023; Alabi et al., 2024a; Zebaze et al., 2024; Zhang et al., 2024a) which include African languages. We identified nineteen relevant papers in total, showing progress but highlighting the need for further work, especially in applying these insights to improve models’ performance.
Ethics, Bias, and Fairness There are also a few works that focus on analyzing biases in NLP models and proposing mitigation strategies to promote fairness. Among the 24 relevant papers identified, gender bias (Costa-jussà et al., 2022) and cultural bias (Shan et al., 2023; Olatunji et al., 2023a; Magomere et al., 2024) were the most commonly addressed. However, much remains to be done to address bias and ensure equity across diverse linguistic and cultural contexts.
Human-Centered NLP Human-centric research in NLP has gained prominence in both the NLP community (Soni et al., 2024) and the field of human-computer interaction (HCI), which has seen a massive uptake on work about LLMs (Pang et al., 2025). However, such work remains scarce in African NLP—for example, only six relevant papers have been presented at CHI, a leading HCI conference, and its African counterpart, AfriCHI.
NLP Applications Our analysis identified 23 relevant papers that focus on the development of NLP technologies aimed at application areas such as education (Corallo and Varde, 2023), health (Abdulhamid et al., 2023), agriculture (Akera et al., 2019), banking (Skiredj et al., 2024), and humanitarian response (Öktem et al., 2021).
@@ 5 Discussion and Future Directions
Scaling beyond the top-10 resourced languages: It is important to invest efforts in supporting African languages beyond Swahili and other widely spoken ones, ensuring that lower-resourced and endangered languages also receive attention in preservation, research, and technological integration. Unlike other regions, Africa has more than 100 languages with at least more than 1M speakers, so scaling beyond the top-10 is urgently needed.
@@
Towards more multi-cultural dataset creation: Our analysis shows that many benchmark datasets created for these languages especially large-scale datasets such as Flores are based on translated texts. Some of these lack cultural contexts of the users of this technology, and can amplify bias towards western concept especially for QA tasks (Romero et al., 2024) or reduce performance due to lack of cultural understanding (Akinade et al., 2023; Yao et al., 2024). There is a need to go beyond scaling by translation to building realistic multicultural datasets for African languages.
@@
Towards more multimodal models for African languages At the moment, there is significantly less research on developing Visual LMs or VLMs for African languages. Existing VLMs perform poorly on African languages and cultural understanding of artifacts from African countries (Nayak et al., 2024; Winata et al., 2024a). This makes it difficult to answer culture-specific questions over images, and also amplifies cultural bias related to Africans identity, lifestyles, occupations, food, clothing, among others (Chiu et al., 2024).
@@
More investment in speech processing tasks: Many languages are often rather spoken than written in Africa, however, only few languages have large-scale datasets covering over 100 hours for ASR. While platforms such as Mozilla CommonVoice have democratized the data collection, only few languages used the platforms due to lack of attribution by the contributors of the datasets (Rajab et al., 2025) and the difficulty of finding volunteers. Aside from popular tasks like ASR, other tasks such as text-to-speech, audio classification, speech translation, and many others lack datasets. To develop NLP models that would be used by many African communities, it must be accompained with a speech component.
@@
More exploration of other generation tasks Most benchmarks for African languages have primarily focused on NLU and MT, partly due to the relative ease of creating datasets for these tasks. However, more attention should be given to language generation tasks—such as text summarization, grammar correction, dialog and table-to-text generation—which are typically more challenging to build datasets for but are equally important for the development of robust language technologies.
@@
More development of African-centric LLMs Current LLMs, especially the open-weight versions, often do not officially include African languages in their pre-training (Aryabumi et al., 2024), and when they do, only few high-resource ones such as Afrikaans and Swahili with more than 1B tokens are covered (Yang et al., 2024). There is need to develop or adapt LLMs to more African languages. This should be accompanied by evaluation datasets that are culturally relevant, sector-specific, knowledge-intensive, and require reasoning to assess LLMs in truly low-resource languages.
@@
Towards human-centered, application-driven NLP Understanding human-centric factors and developing methods in the context of applications is essential to ensure that NLP technologies genuinely benefit users. More research is needed to explore the shared and distinct human-centric needs of African language-speaking communities, both within the continent and in comparison to other global contexts.
@@ 6 Conclusion
Although we used a systematic survey approach that combines automated and expert-driven methods, it is possible that relevant literature focusing on or evaluating African languages within the past five years was missed—particularly if such work was not published in prominent ACL or ML/NLP venues. Similarly, our automated method, which relied on the Semantic Scholar API, may not have correctly indexed all relevant papers from the selected venues. In addition, our manual annotation of 50 examples revealed that our filtering process using GPT-4o on HCI and arXiv papers resulted in some false negatives, including the exclusion of papers on computer vision (though not on NLP). However, a few relevant NLP papers for African languages may have been omitted as well.
@@ Acknowledgments
Jesujoba Alabi was funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) – Project-ID 232722074 – SFB 1102. We are grateful to Aravind Krishnan for their feedback on the manuscript. Finally, we would like to thank OpenAI for providing us access for providing API credits through their Researcher Access API Program.
@@ Appendix A Data collection venue
We collected data from 11 venue categories, as shown in Table 2. While we report both the total number of papers initially retrieved and those ultimately analyzed, not all venues had retrievable papers, and not all contributed to the final analysis (for example, CSCW, UIST). The ACL Summit category had the fewest papers retrieved, while ArXiv, as expected, had the most. Due to the size, and presence of several irrelevant papers from ArXiv and HCI venues, we filtered them using GPT-4o, as described in Appendix B. In addition to the automatically retrieved papers, we also included expert-recommended papers from various venues. Here, experts refers to African researchers who have contributed to African language research over the past five years.
@@
Category Venues #papers #relevant Semantic Scholar AI Conferences AAAI, IJCAI 33 3 ARXIV arXiv.org 643 143 HCI Conferences CHI, AfriCHI, CSCW, UIST 94 6 IR Knowledge CIKM, SIGIR, TREC, WWW 19 7 ML Conferences Neurips, ICML, ICLR 18 5 NLP Conferences ACL, COLING, EACL, EMNLP, LREC, NAACL 192 153 NLP Journals LREA, TACL, TALLIP 19 15 Speech Conference ICASSP, INTERSPEECH, SLT 68 53 Speech Journal TASLP, Speech Communication 10 5 Workshops AfricaNLP, DeepLo, MRL, WOAH, WMT 123 115 ACL Summits EAMT, MT Summit 6 5 Expert Collection Same venues as above, and others e.g. RAIL 230 221 Total 1454 734
@@ Appendix B Data filtering
Due to the large volume of papers retrieved from ArXiv and the presence of many irrelevant ones, we opted for a coarse filtering approach rather than a fine-grained one. Specifically, we filtered papers based on whether they were related to artificial intelligence. To assist with this process, we leveraged GPT-4o. Additionally, we included papers from HCI conferences in this filtering step. To address this, we filtered the papers by prompting GPT-4o777the 2024-08-06 version to classify whether they the papers are AI-related based on their title and abstract (We provide our prompt in LABEL:lst:prompt). We manually evaluated 50 papers from the automatic filtering—25 from each category—and obtained an accuracy of 92%, where the misses are false negatives.
Listing 1: Prompt used for filtering the papers
You are an assistant that determines whether a research paper is related to Artificial Intelligence (AI)."},
Determine if the following paper is related to Artificial Intelligence (AI).
Respond with either AI-related or Not AI-related only.
@@
Appendix C Other tasks researchers work on?
@@
Text with Images or Video: There are a few works on multimodality that combine text with images or video, such as image captioning (Abdulmumin et al., 2022), keyword localization (Olaleye et al., 2022; Nortje et al., 2024), multimodal machine translation (MT) (Abdulmumin et al., 2022; Hatami et al., 2024), text-to-image generation (Leong et al., 2022; Magomere et al., 2024), and visual question answering (Leong et al., 2022). A notable contribution is the development of a multimodal LLM for Amharic (Andersland, 2024). Sign languages, which are inherently multimodal due to their use of visual and spatial cues, have also received attention, with few resource-focused papers addressing sign language recognition (Elhagry and Gla, 2021), sign-to-speech (Kolawole et al., 2022), and sign language translation (Gueuwou et al., 2023b, a).
Language Generation: Our analysis shows that sequence generation tasks—such as text generation (Chang et al., 2020; Ramalepe et al., 2023), text summarization (Ouyang et al., 2019; Zaki et al., 2020; Uthus et al., 2023; Hasan et al., 2021; Pfeiffer et al., 2023), and spelling or grammatical error correction (Adouane et al., 2019; Gezmu et al., 2021a; Keita et al., 2024)—are among the least explored for African languages. Within the text generation category, existing work also includes data-to-text (Gehrmann et al., 2023), question generation (Asai et al., 2024), title or topic generation (Adebara et al., 2024; Meyer et al., 2024), paragraph generation, and automatic diacritization (Olawole et al., 2024; Ojo et al., 2023).

### v3

@@ 4.4 What resources are available?
Given that African languages are low-resource, efforts over the past five years have focused on developing resources—both for individual languages and at a large-scale multilingual level, particularly in the form of datasets. Our analysis shows that 401 papers (about 45%) involve dataset creation, with 126 papers creating datasets via translation. Many large multilingual datasets—such as NLU’s SIB-200 and Taxi-1500—are based on English translations. Similarly, speech datasets like FLEURS, FLEURS-R, and SpeechTaxi rely on translated data, which often lacks cultural relevance and may exhibit translationese (Koppel and Ordan, 2011; Bizzoni et al., 2020), limiting their real-world applicability. Table˜1 lists examples of datasets and models, with a more comprehensive list in Appendix˜E.
@@ 4.5 What are the efficient or low-resource techniques used?
Given the low-resource status of African languages, researchers often rely on specialized techniques. Figure˜7 (Appendix˜D) summarizes those used in the analyzed papers; we highlight six below.
Transfer Learning Our analysis shows that about 49% of the papers used transfer learning, primarily through word embeddings, or by fine-tuning pretrained models or using them as feature extractors. This approach has been effective across many languages, including low-resource ones, largely due to the availability of multilingually pretrained models that support cross-lingual transfer (Pfeiffer et al., 2021). However, a major challenge arises when the target African language is not well represented in the pretrained model or its tokenizer. Common strategies to address this include adaptive pretraining (Pfeiffer et al., 2021; Alabi et al., 2022; Adelani et al., 2024; Meyer et al., 2024), and embedding initialization (Liu et al., 2024; Quinjica and Adelani, 2024; Dobler and de Melo, 2023).
@@
Data Augmentation Data augmentation is another popular low-resource technique, used in 8% of the analyzed papers, and it has been successfully used for various tasks such as language modeling (Adelani et al., 2024; Singh et al., 2024; Azime et al., 2024), machine translation via back translation (Reid et al., 2021; Adelani et al., 2021b), and other tasks (Zhang et al., 2024b).
@@
Other methods Other techniques including annotation projection (AP), meta learning, and multi-task learning (MTL) are described in Appendix˜D.
@@ 4.6 What are the other themes?
Interpretability and Analysis (IA): Methods on interpretability and model analysis have gained increasing attention in NLP in recent years (Mosbach et al., 2024). Although few studies focus exclusively on African languages (Chimoto et al., 2024), there are several multilingual IA works on MT (Ahia et al., 2021; Adebara et al., 2022a; Mohammadshahi et al., 2022b), speech models (Osakuade and King, 2024), and LLM (Ahia et al., 2023; Alabi et al., 2024a; Zebaze et al., 2025; Zhang et al., 2024a; Shafayat et al., 2024) which include African languages. We identified 38 relevant papers in total, showing progress but highlighting the need for further work, especially in applying these insights to improve models’ performance.
Ethics, Bias, and Fairness: There are also a few works that focus on analyzing biases in NLP models and proposing mitigation strategies to promote fairness. Among the 29 relevant papers identified, gender bias (Costa-jussà et al., 2022) and cultural bias (Shan et al., 2023; Olatunji et al., 2023a; Magomere et al., 2025) were the most commonly addressed. However, much remains to be done to address bias and ensure equity across diverse linguistic and cultural contexts.
Human-Centered NLP: Human-centric research in NLP has gained prominence in both the NLP community (Soni et al., 2024) and the field of human-computer interaction (HCI), which has seen a massive uptake on work about LLMs (Pang et al., 2025). However, such work remains scarce in African NLP—for example, only 11 relevant papers have been presented at CHI, a leading HCI conference, and its African counterpart, AfriCHI.
NLP Applications: Our analysis identified 25 relevant papers that focus on the development of NLP technologies aimed at application areas such as education (Corallo and Varde, 2023), health (Abdulhamid et al., 2023), agriculture (Akera et al., 2019), banking (Skiredj et al., 2024), and humanitarian response (Öktem et al., 2021).
@@ 5 Discussion and Future Directions
Scaling beyond the top-10 resourced languages: It is important to invest efforts in supporting African languages beyond Swahili and other widely spoken ones, ensuring that lower-resourced and endangered languages also receive attention in preservation, research, and technological integration. Unlike other regions, Africa has more than 100 languages with more than 1M speakers, so scaling beyond the top-10 is urgently needed.
@@
Towards more multi-cultural dataset creation: Our analysis shows that many benchmark datasets created for these languages, especially large-scale datasets such as Flores, are based on translated texts. Some of these lack cultural contexts of the users of this technology, and can amplify bias towards Western concepts especially for QA tasks (Romero et al., 2024) or reduce performance due to lack of cultural understanding (Akinade et al., 2023; Yao et al., 2024). There is a need to go beyond scaling by translation to building realistic multicultural datasets for African languages.
@@
Towards more multimodal models for African languages: At the moment, there is significantly less research on developing Visual LMs (VLMs) for African languages. Existing VLMs perform poorly on African languages and cultural understanding of artifacts from African countries (Nayak et al., 2024; Winata et al., 2024a). This makes it difficult to answer culture-specific questions over images, and also amplifies cultural bias related to Africans identity, lifestyles, occupations, food, clothing, among others (Chiu et al., 2024).
@@
More investment in speech processing tasks: Many languages in Africa are often spoken rather than written. However, only few languages have large-scale datasets covering over 100 hours for ASR. While platforms such as Mozilla CommonVoice have democratized the data collection, only few languages used the platforms due to lack of attribution by the contributors of the datasets (Rajab et al., 2025) and the difficulty of finding volunteers. Aside from popular tasks like ASR, other tasks such as text-to-speech, audio classification, speech translation, and many others lack datasets. To develop NLP models that would be used by many African communities, it must be accompained with a speech component.
@@
More exploration of other generation tasks: Most benchmarks for African languages have primarily focused on NLU and MT, partly due to the relative ease of creating datasets for these tasks. However, more attention should be given to language generation tasks—such as text summarization, grammar correction, dialog and table-to-text generation—which are typically more challenging to build datasets for but are equally important for the development of robust language technologies.
@@
More development of African-centric LLMs: Current LLMs, especially the open-weight versions, often do not officially include African languages in their pre-training (Aryabumi et al., 2024), and when they do, only few high-resource ones such as Afrikaans and Swahili with more than 1B tokens are covered (Yang et al., 2024). There is need to develop or adapt LLMs to more African languages. This should be accompanied by evaluation datasets that are culturally relevant, sector-specific, knowledge-intensive, and require reasoning to assess LLMs in truly low-resource languages.
@@
Towards human-centered, application-driven NLP: Understanding human-centric factors and developing methods in the context of applications is essential to ensure that NLP technologies genuinely benefit users. More research is needed to explore the shared and distinct human-centric needs of African language-speaking communities, both within the continent and in comparison to other global contexts.
@@ 6 Conclusion
Coverage Limitations:
Although we used a systematic survey approach that combines automated and expert-driven methods, one of the limitations of this survey is that it primarily focused on papers published in top-tier conferences and journals, which may exclude relevant works published within African venues and other venues. While we included papers from AfricaNLP, and AfriCHI, which are both African-centric venues, local research published in other venues may offer additional insights that are not captured in this study. Future work should broaden the scope to incorporate Africa-specific studies to provide more comprehensive understanding of the field and ensure that insights from the different parts of Africa are adequately represented. Similarly, our automated method, which relied on the Semantic Scholar API, may not have correctly indexed all relevant papers from the selected venues.
Potential False Negatives in Data Selection: Our manual annotation of 100 examples revealed that our filtering process using GPT-4o on HCI and arXiv papers resulted in a false negative. It is possible that a few relevant NLP papers on African languages were also omitted, particularly since large language models like GPT-4o are sensitive to prompts, and we only tested a single prompt.
Lack of dataset sheet for some papers: Some papers such as  (Li et al., 2022),  (Nguyen et al., 2024a) and toolkit papers such as  (Emezue and Dossou, 2020a) and (Muite and Kizito, 2022) did not provide the list of languages covered despite covering hundred and thousands of languages. Although we included them in our analysis, we could not provide the studied languages.
Lack of social and ethical dimensions: of African NLP development. This survey focuses primarily on the technical landscape of African NLP, including datasets, tasks, languages, and modeling approaches, and does not systematically address social, cultural, or ethical dimensions such as community involvement, consent, bias, or potential harms of NLP systems. Human-centered evaluations and assessments of societal impact are beyond the scope of this survey. Future research should incorporate ethical, cultural, and social considerations to ensure that African NLP technologies are developed responsibly and equitably.
@@ Acknowledgments
Jesujoba Alabi was funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) – Project-ID 232722074 – SFB 1102. David Adelani acknowledges the funding of IVADO and the Canada First Research Excellence Fund. We are grateful to Aravind Krishnan and Max Rausch-Dupont for their feedback on the manuscript. Finally, we would like to thank OpenAI for providing API credits through their Researcher Access API Program.
@@ Appendix A Data collection venue
We collected data from 11 venue categories, as shown in Table˜2. While we report both the total number of papers initially retrieved and those ultimately analyzed, not all venues had retrievable papers, and not all contributed to the final analysis (for example, CSCW, UIST). The ACL Summit category had the fewest papers retrieved, while ArXiv, as expected, had the most. Due to the size, and presence of several irrelevant papers from ArXiv and HCI venues, we filtered them using GPT-4o, as described in Appendix˜B. In addition to the automatically retrieved papers, we also included expert-recommended papers from various venues. Here, experts refers to African researchers who have contributed to African language research over the past five years.
@@
  32. 32.
    FAccT: ACM Conference on Fairness, Accountability, and Transparency
  33. 33.
    RANLP: International Conference Recent Advances in Natural Language Processing
Category Venues #papers #relevant Semantic Scholar AI Conferences AAAI, IJCAI 2222 6 ARXIV arXiv.org 11680 188 HCI Conferences CHI, AfriCHI, CSCW, UIST 1290 8 IR Knowledge CIKM, SIGIR, TREC, WWW 1507 8 ML Conferences Neurips, ICML, ICLR, FAccT 4804 10 NLP Conferences ACL, COLING, EACL, EMNLP, LREC, NAACL, RANLP 2654 209 NLP Journals LREA, TACL, TALLIP 124 14 Speech Conference ICASSP, INTERSPEECH, SLT 2412 92 Speech Journal TASLP, Speech Communication 279 11 Workshops AfricaNLP, DeepLo, MRL, WOAH, WMT 360 124 ACL Summits EAMT, MT Summit 52 5 Expert Collection Same venues as above, and others e.g. RAIL 209 209 Total 27617 884
@@ Appendix B Data filtering
Due to the large volume of papers retrieved from ArXiv and the presence of many irrelevant ones, we opted for a coarse filtering approach rather than a fine-grained one. Specifically, we filtered papers based on whether they were related to natural language processing. To assist with this process, we used GPT-4o. Additionally, we included papers from HCI conferences in this filtering step. To address this, we filtered the papers by prompting GPT‑4o101010the 2024-08-06 version to classify them into two categories—relevant and not-relevant—based on their title and abstract (our prompt is provided in LABEL:lst:prompt). We then manually evaluated 100 papers from the automatic filtering—50 from each category—and obtained an accuracy of 99%, with all errors being false negatives.
Listing 1: Prompt used for filtering the ArXiv and HCI papers
You are an expert research assistant in Natural Language Processing (NLP).
Given a papers title and abstract, decide whether the paper should be classified as "relevant" or "not relevant".
A paper should be classified as relevant if it is about:
- Natural language processing (NLP) or computational linguistics
- Speech processing (e.g., recognition, synthesis, spoken dialogue)
- Multimodal tasks involving language (e.g., vision-language, speech-text, OCR with text processing, audio-text)
- Dataset curation, annotation, or benchmarks for NLP/multimodal tasks
- Core NLP tasks (e.g., text classification, parsing, entity linking, translation, summarization, information extraction, question answering, dialogue systems, sentiment analysis, semantics, discourse, topic classification, misinformation detection)
- Information retrieval, text mining, or data mining involving text or language understanding
- Evaluation metrics for NLP tasks (e.g., BLEU, ROUGE, COMET, etc.)
- Language modeling
- Language generation
- Ethics, bias, fairness, interpretability, safety in language technology systems
- Human-centered NLP (e.g., user-centered design of language technology, human-LLM interaction, social impact of language technology)
- Applied NLP in specialized domains (e.g., agriculture, biomedical, legal, education, cultural analytics, computational social science, NLP for social good)
Otherwise, the paper should be classified as not relevant.
Output only one label: "relevant" or "not relevant".
@@
Appendix C Other tasks researchers work on
@@
Text with Images or Video: There are a few works on multimodality that combine text with images or video, such as image captioning (Abdulmumin et al., 2022), keyword localization (Olaleye et al., 2022; Nortje et al., 2024), multimodal machine translation (MT) (Abdulmumin et al., 2022; Hatami et al., 2024), text-to-image generation (Leong et al., 2022; Magomere et al., 2025), and visual question answering (Leong et al., 2022). A notable contribution is the development of a multimodal LLM for Amharic (Andersland, 2024). Sign languages, which are inherently multimodal due to their use of visual and spatial cues, have also received attention, with few resource-focused papers addressing sign language recognition (Elhagry and Gla, 2021), sign-to-speech (Kolawole et al., 2022), and sign language translation (Gueuwou et al., 2023b, a).
Language Generation: Our analysis shows that sequence generation tasks—such as text generation (Chang et al., 2020; Ramalepe et al., 2023), text summarization (Ouyang et al., 2019; Zaki et al., 2020; Uthus et al., 2023; Hasan et al., 2021; Pfeiffer et al., 2023), and spelling or grammatical error correction (Adouane et al., 2019; Gezmu et al., 2021a; Keita et al., 2024)—are among the least explored for African languages. Within the text generation category, existing work also includes data-to-text (Gehrmann et al., 2023), question generation (Asai et al., 2024), title or topic generation (Adebara et al., 2024; Meyer et al., 2024), paragraph generation, and automatic diacritization (Olawole et al., 2024; Ojo et al., 2025).